# Thousand Voices of Trauma: A Large-Scale Synthetic Dataset for Modeling Prolonged Exposure Therapy Conversations

### v1

@@ 1 Introduction
The intersection of mental health care and artificial intelligence presents unprecedented opportunities alongside significant challenges. AI system development faces particular obstacles in trauma-focused therapy due to the sensitive nature of patient experiences and strict privacy regulations, which make the collection of real-world data extremely challenging (Mathur et al., 2022). Moreover, existing datasets frequently lack the diversity and clinical depth needed to train robust AI systems capable of serving diverse populations effectively Marr (2023). Prolonged Exposure (PE) therapy—an evidence-based treatment for post-traumatic stress disorder (PTSD) (Foa et al., 2007)—offers a structured therapeutic approach that could especially benefit from AI support. However, current mental health conversation datasets are often too small (Zantvoort et al., 2024), lack demographic diversity (Sadeh-Sharvit, 2024), and do not capture the nuanced progression of trauma-focused therapy sessions (Mathur et al., 2022).
@@
This dataset also addresses other real-world limitations that often hinder mental health research and safe AI model development. For example, while privacy concerns often restrict data access, synthetic data can circumvent typical ethical and legal barriers. It also overcomes other common issues with real-world data, such as incompleteness, inconsistency, and small sample sizes—especially among minority groups. By balancing representation across diverse populations, trauma types, and racial or ethnic minorities, it helps mitigate inherent biases. For instance, the NIMH reported in 2021 that 14.5 million U.S. adults (5.7%) experienced severe major depressive episodes, with higher rates among females (10.3%) than males (6.2%) and the highest prevalence among those aged 18–25 (18.6%). Synthetic data can compensate for such imbalances, enhancing model training and analysis.
@@ 2 Related Work
Synthetic datasets show promise in mental health applications, with studies exploring LLM-based data generation to address data scarcity. Wu et al. (Wu et al., 2023, 2024) introduced zero-shot and few-shot learning frameworks to augment PTSD diagnostic datasets, producing synthetic transcripts that outperform baselines. The latter work used role-prompting and structured prompts to create realistic synthetic clinical interviews. However, these approaches have not been specifically adapted to the structured narratives essential for PE therapy.
Efforts to enhance demographic diversity in synthetic datasets are growing. Mori et al. (2024) and Lozoya et al. (2023) examined how synthetic data reflects demographic variation, highlighting biases in LLM outputs, especially regarding race and gender, and stressing fairness in mental health datasets. Techniques like patient vignette simulation (Denecke, 2024) and adaptive prompts for non-English contexts (e.g., SAPE for Spanish (Lozoya et al., 2024)) show early progress toward inclusivity. However, trauma-type diversity and PE-specific scenarios remain unexplored. Additionally, Chen et al. (2024) underscored the need for systematic benchmarking using metrics like F1-score, AUC, and balanced accuracy, but these tools have not yet been applied to datasets focused on diverse trauma types or PE therapy.
@@
While synthetic dataset generation using LLMs (Wu et al., 2023, 2024; Xu et al., 2024) has advanced in addressing general clinical challenges, gaps in the literature remain for PE therapy. These include the lack of trauma-type diversity, limited demographic inclusivity evaluation, insufficient alignment with frameworks like DSM-5 PTSD criteria, and underdeveloped ethical safeguards specific to trauma-focused contexts. Synthetic datasets can also help mitigate representation biases in AI models for mental healthcare delivery. For example, American Psychiatric Association (2013) reported that non-Hispanic White adults (25.0%) were more likely to receive mental health services than non-Hispanic Black (18.3%), Hispanic (17.3%), and Asian (13.9%) adults. Including such underrepresented groups in synthetic datasets might partially address the training data gap. This paper seeks to bridge these gaps by exploring the existing knowledge base and identifying pathways to tailor synthetic data generation for PE therapy applications.
@@ 3.6 Dataset Statistics
The dataset comprises 500 simulated participants (ages 18-80 years, M = 49.3). The gender distribution includes 247 male (49.4%), 222 female (44.4%), and 31 non-binary (6.2%) participants. The ethnicity distribution is as follows: Latin American (70, 14.0%), North American (67, 13.4%), South East Asian (64, 12.8%), South Asian (63, 12.6%), European (63, 12.6%), Oceania (62, 12.4%), Middle Eastern (59, 11.8%), and African (52, 10.4%). Regarding relationship status, participants were predominantly married (180, 36.0%) or single (149, 29.8%), with others reporting being in a relationship (70, 14.0%), divorced (51, 10.2%), or widowed (50, 10.0%). See Figure 1 & 2 for more details.
@@ 4 Expert Evaluation of Synthetic PE Therapy Sessions
The evaluation yielded valuable, multifaceted feedback. Experts consistently recognized the dataset’s strengths in capturing detailed patient narratives, with ratings ranging from “Somewhat detailed” (n=3) to “Very detailed” (n=4). Furthermore, the simulated patient’s engagement was rated positively, ranging from “Moderately” engaged (n=4) to “Extremely” engaged (n=3), indicating the model’s success in generating realistic and emotionally resonant patient responses crucial for PE simulations. Identified strengths included vivid trauma descriptions, significant emotional depth, and the portrayal of realistic novice therapist approaches – elements particularly useful for training applications.
The feedback also highlighted areas reflecting the inherent challenges of synthetically generating complex therapeutic interactions. Ratings for perceived value (ranging from “Not valuable” (n=2) to “Valuable” (n=1)) and appropriateness (from “Not appropriate” (n=3) to “Somewhat appropriate” (n=4)) varied, underscoring the difficulty in fully replicating nuanced clinical judgment and meeting all expert expectations synthetically. Similarly, the AI therapist’s perceived skill level (rated between “Novice” (n=2) and “Competent” (n=3)) suggests the current simulation captures less experienced therapist behavior more readily than expert-level interactions.
@@ 5.2 Benchmark Design
The expected emotional trajectory follows a recognizable pattern (Cowden Hindash et al., 2020; Bluett et al., 2014; Center for the Treatment and Study of Anxiety, 2023): initial anxiety during orientation, peaking distress during imaginal exposure, followed by gradual reduction through reinforcement and processing (see Figure 6). SUDS (Subjective Units of Distress Scale) are used to measure self-reported distress during exposure. Although "Monitoring SUDS Ratings" is a distinct segment in our dataset, it occurs concurrently with the "Imaginal Exposure" phase in the PE protocol. Therefore, their emotional intensity ratings are expected to align closely, reflecting shared peak distress dynamics.
@@ 5.3 Evaluation Metrics
  • Sequence Similarity (Pearson Correlation): Measures the linear correlation between the model’s predicted emotional intensity sequence and the baseline sequence across the six phases. It assesses alignment in both magnitude and the relative ranking of distress levels.
@@ 5.4 Baseline Establishment and Validation
To establish a reference for comparison, we generated a Baseline Trajectory using zero-shot emotional intensity ratings from Claude Sonnet 3.5. Models assign scores on a 1 (calm) to 10 (extreme distress) scale for each of the six conversation segments per client profile. Claude Sonnet 3.5 was chosen for its strong zero-shot natural language understanding capabilities, suggesting suitability for interpreting emotional nuances without task-specific training.
As shown in Figure 6, the resulting Baseline Trajectory aligns with established PE therapy patterns: peak distress occurs during the Imaginal Exposure and concurrent Monitoring SUDS Ratings phases, with gradual reductions during subsequent Reinforcing, Eliciting Thoughts/Feelings, and Processing phases. This correlates with the human evaluation and confirms the baseline’s suitability for evaluating future models.
@@ 5.6 Results and Discussion
Interestingly, Mistral Small ($S_{\text{abs}}=0.59$) showed the weakest alignment, performing worse than the smaller Mistral 7B Instruct model ($S_{\text{abs}}=0.63$). This finding highlights the importance of instruction-following capabilities for the benchmark. Mistral 7B Instruct is specifically tuned for following commands, crucial for adhering to the benchmark’s requirements (e.g., correct scoring scale and format). Mistral Small’s observed difficulties generating correctly formatted responses support the hypothesis that its weaker performance stems from poorer instruction adherence or task suitability rather than the model size.
Overall, these comparisons validate the benchmark’s effectiveness in providing quantitative measures of alignment and differentiating between models with varying capabilities for tracking emotional intensity in simulated PE therapy conversations.
@@ 6 Future Directions
Building on this work, future research should leverage synthetic therapy session conversations to augment real-world datasets, enhancing AI model training with a focus on clinical fidelity. This dataset can serve as a novel tool for therapist education. For example, simulated patient profiles can be used in role-playing, case studies, and interactive simulations, which can lead to more effective therapist training.
This dataset can also help develop supportive technologies for patients. For example, this dataset can potentially be extended to facilitate supportive chatbot development. Chatbots trained on nuanced emotional and therapeutic data can expand access to mental health support, offering anonymous, always-available assistance, especially for those facing barriers like cost, location, or stigma. While not a replacement for human therapists, they can serve as a first point of contact, support mild to moderate concerns, and detect crises, directing users to emergency services.
@@ 8 Ethical Considerations
It is important to emphasize that this dataset is entirely synthetic and does not involve any real individuals or their personal experiences. The data generation process was designed to capture potential patterns and relationships within the context of therapy but should not be interpreted as a direct reflection of any specific population or individual’s lived experiences.

### v2

@@ 1 Introduction
The intersection of mental health care and artificial intelligence presents unprecedented opportunities alongside significant challenges. AI system development faces particular obstacles in trauma-focused therapy due to the sensitive nature of patient experiences and strict privacy regulations, which make the collection of real-world data extremely challenging (Mathur et al., 2022).
Moreover, existing datasets frequently lack the diversity and clinical depth needed to train robust AI systems capable of serving diverse populations effectively Marr (2023). Prolonged Exposure (PE) therapy—an evidence-based treatment for post-traumatic stress disorder (PTSD) (Foa et al., 2007)—offers a structured therapeutic approach that could especially benefit from AI support.
However, current mental health conversation datasets are often too small (Zantvoort et al., 2024), lack demographic diversity (Sadeh-Sharvit, 2024), and do not capture the nuanced progression of trauma-focused therapy sessions (Mathur et al., 2022).
@@
This dataset also addresses other real-world limitations that often hinder mental health research and safe AI model development. For example, while privacy concerns often restrict data access, synthetic data can circumvent typical ethical and legal barriers. It also overcomes other common issues with real-world data, such as incompleteness, inconsistency, and small sample sizes—especially among minority groups.
By balancing representation across diverse populations, trauma types, and racial or ethnic minorities, the dataset helps mitigate inherent biases. For instance, the NIMH reported in 2021 that 14.5 million U.S. adults (5.7%) experienced severe major depressive episodes, with higher rates among females (10.3%) than males (6.2%) and the highest prevalence among those aged 18–25 (18.6%). Synthetic data can compensate for such imbalances, enhancing model training and analysis.
@@ 2 Related Work
Synthetic datasets show promise in mental health applications, with studies exploring LLM-based data generation to address data scarcity. Wu et al. (Wu et al., 2023, 2024) introduced zero-shot and few-shot learning frameworks to augment PTSD diagnostic datasets, producing synthetic transcripts that outperform baselines. The latter work used role-prompting and structured prompts to create realistic synthetic clinical interviews.
However, these approaches have not been specifically adapted to the structured narratives essential for PE therapy. Efforts to enhance demographic diversity in synthetic datasets are growing. Mori et al. (2024) and Lozoya et al. (2023) examined how synthetic data reflects demographic variation, highlighting biases in LLM outputs, especially regarding race and gender, and stressing fairness in mental health datasets. Techniques like patient vignette simulation (Denecke, 2024) and adaptive prompts for non-English contexts (e.g., SAPE for Spanish (Lozoya et al., 2024)) show early progress toward inclusivity. However, trauma-type diversity and PE-specific scenarios remain unexplored. Additionally, Chen et al. (2024) underscored the need for systematic benchmarking using metrics like F1-score, AUC, and balanced accuracy, but these tools have not yet been applied to datasets focused on diverse trauma types or PE therapy.
@@
While synthetic dataset generation using LLMs (Wu et al., 2023, 2024; Xu et al., 2024) has advanced in addressing general clinical challenges, gaps in the literature remain for PE therapy. These include the lack of trauma-type diversity, limited demographic inclusivity evaluation, insufficient alignment with frameworks like DSM-5 PTSD criteria, and underdeveloped ethical safeguards specific to trauma-focused contexts. Synthetic datasets can also help mitigate representation biases in AI models for mental healthcare delivery.
For example, American Psychiatric Association (2013) reported that non-Hispanic White adults (25.0%) were more likely to receive mental health services than non-Hispanic Black (18.3%), Hispanic (17.3%), and Asian (13.9%) adults. Including such underrepresented groups in synthetic datasets might partially address the training data gap. This paper seeks to bridge these gaps by exploring the existing knowledge base and identifying pathways to tailor synthetic data generation for PE therapy applications.
@@ 3.6 Dataset Statistics
The dataset comprises 500 simulated participants (ages 18-80 years, M = 49.3). The gender distribution includes 247 male (49.4%), 222 female (44.4%), and 31 non-binary (6.2%) participants.
The ethnicity distribution is as follows: Latin American (70, 14.0%), North American (67, 13.4%), South East Asian (64, 12.8%), South Asian (63, 12.6%), European (63, 12.6%), Oceania (62, 12.4%), Middle Eastern (59, 11.8%), and African (52, 10.4%). Regarding relationship status, participants were predominantly married (180, 36.0%) or single (149, 29.8%), with others reporting being in a relationship (70, 14.0%), divorced (51, 10.2%), or widowed (50, 10.0%). See Figure 1 & 2 for more details.
@@ 4 Expert Evaluation of Synthetic PE Therapy Sessions
The evaluation yielded valuable, multifaceted feedback. Experts consistently recognized the dataset’s strengths in capturing detailed patient narratives, with ratings ranging from “Somewhat detailed” (n=3) to “Very detailed” (n=4). Furthermore, the simulated patient’s engagement was rated positively, ranging from “Moderately” engaged (n=4) to “Extremely” engaged (n=3), indicating the model’s success in generating realistic and emotionally resonant patient responses crucial for PE simulations. Identified strengths included vivid trauma descriptions, significant emotional depth, and the portrayal of realistic novice therapist approaches – elements particularly useful for training applications. The feedback also highlighted areas reflecting the inherent challenges of synthetically generating complex therapeutic interactions.
Ratings for perceived value (ranging from “Not valuable” (n=2) to “Valuable” (n=1)) and appropriateness (from “Not appropriate” (n=3) to “Somewhat appropriate” (n=4)) varied, underscoring the difficulty in fully replicating nuanced clinical judgment and meeting all expert expectations synthetically. Similarly, the AI therapist’s perceived skill level (rated between “Novice” (n=2) and “Competent” (n=3)) suggests the current simulation captures less experienced therapist behavior more readily than expert-level interactions.
@@ 5.2 Benchmark Design
The expected emotional trajectory follows a recognizable pattern (Cowden Hindash et al., 2020; Bluett et al., 2014; Center for the Treatment and Study of Anxiety, 2023): initial anxiety during orientation, peaking distress during imaginal exposure, followed by gradual reduction through reinforcement and processing (see Figure 6). SUDS (Subjective Units of Distress Scale) are used to measure self-reported distress during exposure.
Although "Monitoring SUDS Ratings" is a distinct segment in our dataset, it occurs concurrently with the "Imaginal Exposure" phase in the PE protocol. Therefore, their emotional intensity ratings are expected to align closely, reflecting shared peak distress dynamics.
@@ 5.3 Evaluation Metrics
  • Sequence Similarity (Pearson Correlation): Measures the linear correlation between the model’s predicted emotional intensity sequence and the baseline sequence across the six phases. It assesses alignment in both magnitude and relative ranking of distress levels.
@@ 5.4 Baseline Establishment and Validation
To establish a reference for comparison, we generated a Baseline Trajectory using zero-shot emotional intensity ratings from Claude Sonnet 3.5. Models assign scores on a 1 (calm) to 10 (extreme distress) scale for each of the six conversation segments per client profile. Claude Sonnet 3.5 was chosen for its strong zero-shot natural language understanding capabilities, suggesting suitability for interpreting emotional nuances without task-specific training. As shown in Figure 6, the resulting Baseline Trajectory aligns with established PE therapy patterns: peak distress occurs during the Imaginal Exposure and concurrent Monitoring SUDS Ratings phases, with gradual reductions during subsequent Reinforcing, Eliciting Thoughts/Feelings, and Processing phases. This correlates with the human evaluation and confirms the baseline’s suitability for evaluating future models.
@@ 5.6 Results and Discussion
Interestingly, Mistral Small ($S_{\text{abs}}=0.59$) showed the weakest alignment, performing worse than the smaller Mistral 7B Instruct model ($S_{\text{abs}}=0.63$). This finding highlights the importance of instruction-following capabilities for the benchmark. Mistral 7B Instruct is specifically tuned for following commands, crucial for adhering to the benchmark’s requirements (e.g., correct scoring scale and format). Mistral Small’s observed difficulties generating correctly formatted responses support the hypothesis that its weaker performance stems from poorer instruction adherence or task suitability rather than the model size. Overall, these comparisons validate the benchmark’s effectiveness in providing quantitative measures of alignment and differentiating between models with varying capabilities for tracking emotional intensity in simulated PE therapy conversations.
@@ 6 Future Directions
Building on this work, future research should leverage synthetic therapy session conversations to augment real-world datasets, enhancing AI model training with a focus on clinical fidelity. This dataset can serve as a novel tool for therapist education. For example, simulated patient profiles can be used in role-playing, case studies, and interactive simulations, which can lead to more effective therapist training. This dataset can also help develop supportive technologies for patients. For example, this dataset can potentially be extended to facilitate supportive chatbot development.
Chatbots trained on nuanced emotional and therapeutic data can expand access to mental health support, offering anonymous, always-available assistance, especially for those facing barriers like cost, location, or stigma. While not a replacement for human therapists, they can serve as a first point of contact, support mild to moderate concerns, and detect crises, directing users to emergency services.
@@ 8 Ethical Considerations
This dataset is entirely synthetic and does not involve any real individuals or their personal experiences. While it captures potential patterns within therapy conversations, it should not be interpreted as reflecting any specific population or lived experience.
Our work simulates Prolonged Exposure (PE) therapy for PTSD—a treatment involving sensitive and traumatic content. We recognize the ethical implications related to representation, misuse, and potential harm.
To address these concerns: (1) no real patient data was used—the synthetic dialogues were generated using language models prompted with protocol-adherent scenarios from publicly available PE manuals; (2) we avoid simulating graphic trauma details, focusing on structural and therapeutic fidelity; and (3) any real transcripts used for comparison were collected under IRB-approved protocols with participant consent and de-identified prior to analysis. The dataset release will include usage guidelines and license restrictions to discourage misuse in non-clinical or consumer-facing applications. Our intent is to support clinical NLP research in a privacy-preserving, scalable way—not to replace trained therapists.
We emphasize the need for continued community discussion on the ethical use of synthetic therapy data. Our work does not validate the clinical effectiveness of PE or its synthetic representation. Rather, it investigates whether LLMs can replicate the structure of therapy dialogues for research and training purposes—not simulate real trauma responses or therapeutic outcomes.
@@ 9 Limitations
10 Acknowledgement
This work is supported by the National Science Foundation (NSF) under Grant No. 2326144. Any opinions, findings, & conclusions or recommendations expressed in this work are those of the author(s) and do not necessarily reflect the views of the NSF.